In a recent survey of 850 students, it was found that 476 of them have laptops. a. Find the sample proportion p ̂. b. Find a 95% confidence interval for the actual proportion of students who have laptops. 


1
Expert's answer
2021-04-27T01:02:05-0400

We have sample of 850 students.

Out of which 476 have laptops.

a.)Sample proportion "p = \dfrac{476}{850}"


b.) Confidence interval can be given as "p \pm z\sqrt{\dfrac{p(1-p)}{n}}"


Value of z at 95% confidence level is 1.96


Hence, Confidence interval "= \dfrac{476}{850} \pm 1.96 \sqrt{\dfrac{0.56\times 0.44}{850}}"

"= 0.56 \pm 0.03"